OpenStudy (anonymous):

For which system of equations is (5, 3) the solution? A. 3x – 2y = 9 3x + 2y = 14 B. x – y = –2 4x – 3y = 11 C. –2x – y = –13 x + 2y = –11 D. 2x – y = 7 2x + 7y = 31

OpenStudy (ajprincess):

plug in 5 for x and 3 for y in the equations and see if the left side and right side are equal. if they are equal then (5,3) is a solution.

OpenStudy (anonymous):

That didnt really help much.

OpenStudy (ajprincess):

oh then let me show u an example. let;s consider the first choice. 3x – 2y = 9 3x + 2y = 14 nw consider the lhs of first equation which is 3x-2y. substitutinx 5 for x and 3 for y gives 3*5-2*3=15-6=9 so lhs=rhs so (5,3) is a solution of this equation. consider the lhs of the second equation 3x+2y=3*5+2*3=15+6=21 lhs is not equal to rhs. hence it is not a solution of this equation. hence (5,3) is not a solution of this system of equations. hence this is not the right option